# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v5

@@ Abstract
Efficiently learning and executing long-horizon mobile manipulation (MoMa) tasks is crucial for advancing robotics in household and workplace settings. However, current MoMa models are data-inefficient, underscoring the need for improved models that require realistic-sized benchmarks to evaluate their efficiency, which do not exist. To address this, we introduce the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities), which evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. The benchmark includes 571 human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We benchmark several models, including learning-based models and a neuro-symbolic modular approach combining foundation models with task and motion planning. Learning-based models show suboptimal success rates, even when leveraging pretrained weights, underscoring significant data inefficiencies. However, the neuro-symbolic approach performs significantly better while being more data efficient. Findings highlight the need for more data-efficient learning-based MoMa approaches. addresses this gap by serving as a key benchmark for evaluating the data efficiency of those future models in handling household robotics tasks.
[…]
figureOverview. LAMBDA () is a benchmark comprised of 571 language-conditioned long-horizon (room-to-room and floor-to-floor navigation) mobile manipulation (pick-and-place) tasks paired with human-collected expert demonstrations from both simulated and real-world environments. The illustrated examples in both the simulated and real-world environments span multiple rooms, which is a skill expected of daily household and workspace robots. We use a quadruped robot for its ability to recover from falls and to navigate complex terrain such as stairs, which is essential for floor-to-floor tasks.
@@ I INTRODUCTION
Improving efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Although significant strides have been made in tabletop manipulation with various architectures and approaches for more data-efficient models [1, 2, 3], existing MoMa models rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6]. Collecting such large-scale datasets is resource-intensive and requires substantial time, manpower, and infrastructure.This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse task requirements, and the need for generalization in home and workplace settings. Addressing these challenges demands the development of more data-efficient approaches for MoMa, supported by benchmarks that can guide and evaluate this progress effectively.
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, for long-horizon, language-conditioned MoMa tasks, leaving a significant gap in evaluating and advancing progress in this area. For instance, most lack language instruction annotations for the tasks [7, 8], even though natural language provides an intuitive and accessible interface for specifying robot objectives, especially for household and workplace users. Also, most benchmarks do not support long-horizon tasks [4, 9] requiring navigation across rooms and floors, which are essential for realistic household scenarios. Furthermore, while some benchmarks support simulated environments, they often lack real-world data [10, 11], which are crucial for ensuring models’ effectiveness in practical deployments. Some current benchmark datasets also rely on planner-generated trajectories [12, 13], which lack the natural variability and expertise of human-collected demonstrations. Finally, many benchmarks lack quadruped data [14, 15], which are crucial for terrains like stairs for floor-to-floor navigation, and maintaining stability during tasks.
@@
To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. The dataset consists of 571 high-quality trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in provide natural variability and expertise, better capturing real-world robot operations. Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Its simulated environments enable diverse testing scenarios and object variability, ensuring flexibility and standardization, while real-world data ensures models are prepared for practical deployment. Finally, the inclusion of armed quadrupeds enables evaluation on complex terrains, such as stairs, highlighting their stability and adaptability in challenging conditions.
@@
We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the dataset’s ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future. These results demonstrate ’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved ones for indoor robotics.
Our contributions are summarized as follows:
benchmark: We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings.
High-quality dataset: ’s tasks are paired with 571 human-collected trajectories, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation.
Comprehensive evaluation: We benchmark several models, including BC and neuro-symbolic modular approaches, providing a baseline for future model development for MoMa tasks of this nature.
@@
| Benchmark | Natural Language | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
| SafeAgentBench | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| HumanoidBench | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ |
| Mobile ALOHA | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| BEHAVIOR-1K | ✗ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ |
| ManiSkill-HAB | ✗ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| InfiniteWorld | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ |
| RoboSuite | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✓ |
| RoboCasa | Some | ✗ | ✓ | Some | ✓ | ✓ | ✓ |
| BRMData | ✗ | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| DeMoBot | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ |
| MoMaRT | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
| M3Bench | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| BiGym | ✗ | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ |
| RT-1 | ✓ | ✗ | ✓ | ✓ | ✓ | ✓ | ✗ |
@@ II Related Work
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [7, 8, 16, 13, 17, 18, 19, 15, 9], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications.
@@
Moreover, almost all benchmarks [7, 8, 13, 17, 14, 18, 19, 15, 12, 9, 4], do not support long-horizon tasks involving navigation across rooms and floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. While most benchmarks provide simulated environments, Mobile ALOHA [8] and BRMData [18] lack simulation data entirely. Simulated environments offer diverse scenes and objects for flexible testing across varied scenarios and provide a standardized, portable testbed that can be easily used by others due to their fixed conditions. On the other hand, most lack real-world data [10, 7, 13, 11, 17, 15, 12, 9], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
@@
Most benchmarks provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [10], HumanoidBench [7], BEHAVIOR-1K [16], and InfiniteWorld [11], lack this attribute, limiting their utility for model training and evaluation. M3Bench [12] and ManiSkill-HAB [13] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
@@
Finally, most benchmarks from Table I, with the exceptions of RoboCasa [14], RoboSuite [17], and DeMoBot [19], do not feature armed quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in . Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
@@ III Benchmark
We present LAMBDA (), a benchmark designed to evaluate model data efficiency on long-horizon MoMa tasks. features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings. The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
@@ III-A Challenges of Tasks
Data Efficiency in MoMa. is designed to test the data efficiency of learning-based models by offering a dataset of moderate size, deliberately curated to strike a balance between realism and practicality. While recent data-efficient tabletop manipulation models [1, 2, 3] often rely on double-digit demonstrations, we provide a few hundred demonstrations, recognizing the increased complexity of MoMa. MoMa encompasses not only higher-dimensional observation, state, and action spaces but also significantly larger and more varied task spaces. This deliberate dataset size ensures sufficient diversity for learning while maintaining a manageable scale that avoids the excessive computational and collection costs of unrealistically large datasets.
Complex Spaces. The inclusion of MoMa tasks introduces significant complexity due to the interplay of action, observation, state, and task spaces. Unlike fixed single-arm setups, MoMa requires navigating and interacting within dynamic three-dimensional environments, often spanning multiple rooms or floors. This necessitates the seamless integration of high-level planning and low-level motion control, creating challenges for robotic models to operate effectively. For example, consider a household task where the agent must navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this straightforward pick-and-place task, the complexity of the action, state, and observation spaces introduces substantial variability. The agent might approach the table from different directions, choose from various grasping points on the plate, or follow alternative routes to the destination. This multi-modal nature means that the same task can be solved in multiple ways, further increasing the demands on models to generalize across diverse strategies and environmental configurations. Moreover, the complexity of these integrated spaces enables a broader range of tasks, from transporting objects across floors to rearranging items in dynamic environments. For our benchmark, we focus on pick-and-place tasks, as they represent a significant portion of daily activities and provide a foundational challenge for MoMa models to address.
Long-Horizon Tasks. A key challenge in lies in its dedicated emphasis on long-horizon tasks, where navigation spans multiple rooms and floors in indoor environments such as households and workplaces. These tasks inherently demand long-term planning, often requiring hierarchical strategies that integrate high-level task objectives with low-level motion execution. The agent must account for spatial relationships, sequential task dependencies, and the compounding effects of errors over time. Error propagation is a critical challenge across many learning paradigms, exemplified by covariate shift in BC. For instance, navigating to a distant room to retrieve an object and returning it to another location involves traversing obstacles and staircases while adhering to task objectives and maintaining situational awareness. Such tasks are representative of real-world scenarios where robots are expected to perform complex, multi-step instructions over extended periods.
Realistic Task Distribution. To ensure the tasks in reflect practical needs, we used crowdsourcing to gather a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. This method avoided templated task specifications by capturing natural, human-generated instructions. The resulting dataset encompasses a variety of pick-and-place MoMa tasks, aligning with real-world expectations and exposing models to inherently varied and context-dependent scenarios.

### v6

@@ Abstract
Learning to execute long-horizon mobile manipulation tasks is crucial for advancing robotics in household and workplace settings. However, current approaches are typically data-inefficient, underscoring the need for improved models that require realistically sized benchmarks to evaluate their efficiency. To address this, we introduce the LAMBDA () benchmark111lambdabenchmark.github.ioââLong-horizon Actions for Mobile-manipulation Benchmarking of Directed Activitiesââwhich evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. Our benchmark includes human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We leverage LAMBDA to benchmark current end-to-end learning methods and a modular neuro-symbolic approaches that combines foundation models with task and motion planning. We find that end-to-end methodsâeven when pretrainedâyield lower success rates, while neuro-symbolic methods perform significantly better and require less data.
@@ I INTRODUCTION
Improving data efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Significant strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3], however existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6]. Collecting such large-scale datasets is resource-intensive, requiring substantial time, labor, and infrastructure. This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse tasks, and the need for generalization across environments.
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, leaving a significant gap in evaluating progress in this area. A few works  [7, 8, 9, 10, 4, 11] pair task demonstrations with language instruction annotations and support long-horizon tasks [7, 12, 8, 11], providing an intuitive and accessible interface for specifying robot objectives. Even fewer [9, 11] provide natural language instructions in an open-ended, flexible format rather than relying on fixed templates. Many benchmarks lack real-world data [13, 12, 9, 14, 15, 4] or rely on planner-generated trajectories [16, 10, 17], limiting their practicality. Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide quadruped data.
@@
To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. Our dataset consists of trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in provide natural variability and expertise, better capturing real-world robot operations. Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Simulated environments enable diverse testing scenarios and object variability, while real-world data ensures models are prepared for practical deployment. We collected this dataset with a quadruped robot [18] due to its optimal form-factor suitable for navigating complex terrain.
@@
We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the datasetâs ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
Overall our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings. (2) High-quality dataset:âs tasks are paired with human-collected robot demonstrations, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation. (3) Comprehensive evaluation:We benchmark several models, including BC and modular neuro-symbolic approaches, providing a baseline for future model development for MoMa tasks.
@@
| Benchmark | NL | FF NL | Long-Horizon | Demos | Human-Collected | Sim | Real | Quadruped |
| SafeAgentBench [7] | â | â | â | â | â | â | â | â |
| HumanoidBench [20] | â | â | â | â | â | â | â | â |
| Mobile ALOHAÂ [13] | â | â | â | â | â | â | â | â |
| BEHAVIOR-1KÂ [12] | â | â | â | â | â | â | â | â |
| ManiSkill-HABÂ [16] | â | â | â | â | â | â | â | â |
| InfiniteWorld [8] | â | â | â | â | â | â | â | â |
| M2Diffuser [17] | â | â | â | â | â | â | â | â |
| RoboSuite [19] | â | â | â | â | â | â | â | â |
| BRMData [14] | â | â | â | â | â | â | â | â |
| RoboCasa [9] | Some | â | â | â | Some | â | â | â |
| DeMoBot [15] | â | â | â | â | â | â | â | â |
| MoMaRTÂ [21] | â | â | â | â | â | â | â | â |
| M3Bench [10] | â | â | â | â | â | â | â | â |
| PARTNRÂ [11] | â | â | â | â | â | â | â | â |
| BiGym [22] | â | â | â | â | â | â | â | â |
| RT-1Â [4] | â | â | â | â | â | â | â | â |
| â | â | â | â | â | â | â | â |
@@ II Related Work
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [20, 13, 12, 16, 19, 14, 15, 21, 22], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications. Even fewer benchmarks, such as RoboCasa [9] and PARTNR [11], offer natural language instructions in a free-form, unconstrained format rather than predefined templates. This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in-the-wild.
@@
Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17], do not support long-horizon tasks involving navigation across rooms and floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. As shown in Table I, Mobile ALOHA [13] and BRMData [14] lack simulation data entirely. Simulated environments offer diverse scenes and objects, making it easier to test models across a wide range of scenarios. Their standardized and controlled conditions create a portable testbed that others can readily use, ensuring consistency and reproducibility in evaluations. On the other hand, most lack real-world data [7, 20, 16, 8, 19, 21, 10, 22, 11, 17], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
@@
The majority of benchmarks in Table I provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11] lack this attribute, limiting their utility for model training and evaluation. M3Bench [10], M2Diffuser [17], and ManiSkill-HAB [16] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
@@
Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in . Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
@@ III Benchmark
The LAMBDA () benchmark is designed to evaluate model data efficiency on long-horizon MoMa tasks. features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings. The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
@@ III-A Challenges of Tasks
Data Efficiency in MoMa. We define data efficiency as the ability to achieve relatively strong performance with minimal robot trajectory demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every time step. Non-end-to-end systems are still considered within our definition of data efficiency. For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstration trajectories, are not excluded on this basis. Unlike very large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
While previous tabletop manipulation methods depended on thousands of demonstrations [23], recent models [1, 2, 3] are more data efficient, requiring only tens of demonstrations. In a similar direction, recognizing the increased complexity of MoMa, we provide hundreds of demonstrations for model development and benchmarking. This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient, avoiding the excessive resource costs of unrealistically large datasets.
Long Horizon Tasks in Complex Spaces. MoMa tasks introduce significant complexity by involving larger and more intricate action, observation, state, and task spaces compared to tabletop manipulation. Operating in dynamic three-dimensional environments that span multiple rooms and floors (shown in Figure LABEL:fig:main_splash), these tasks require the seamless integration of high-level planning with low-level motion control. For example, an agent may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe agent might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, sequential dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with behavioral cloning approaches [26]. Unlike fixed single-arm setups, MoMaâs expansive spaces force the agent to contend with a higher degree of variability and uncertainty. Consequently, our benchmark focuses on pick-and-place tasks, such as the ones presented in Figure 1, which serve as a representative foundation for addressing the broader demands of MoMa models in complex, real-world environments.
Crowdsourced Free-Form Natural Language Instructions. Understanding language is crucial for robots in human environments, as it is the most natural way for humans to communicate task specifications. Thus, âs tasks are language-conditioned, meaning goals are not specified via images, state information, or symbolic representations. Language adds complexity, requiring models to interpret and ground linguistic expressions [27] instead of relying on explicit visual or spatial cues.
To ensure tasks in reflect real-world needs, we collected free-form human-generated instructions via crowdsourcing. This provides a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. Unlike system-generated templated commands, âs free-form instructions, seen in Figures LABEL:fig:main_splash & 1, vary in phrasing, detail, and structure, requiring models to generalize across a variety of expressions rather than memorizing fixed patterns. This pushes models to interpret and reason about task objectives. Additionally, these instructions introduce ambiguities and implicit references further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.